OpenStudy (anonymous):

Find the area between the x − axis and f(x) = 1 + 1/x over the interval [1, 10]. Round your answer to five decimal places.

OpenStudy (helder_edwin):

u can use Riemann's sum just like in the previous problem u posted. though you'll have to use a patition of say 100 points

OpenStudy (anonymous):

Really? Riemann sums? Is there a reason he would want to do that? That's kind of ridiculous.

OpenStudy (anonymous):

isn't this just ln(10)?

OpenStudy (anonymous):

with n=100 ?

OpenStudy (anonymous):

The integral of 1/x is ln x, evaluated between 1 and 10 is ln(10). No sums.

OpenStudy (anonymous):

 \[\Large \int\limits_{1}^{10}1+\frac{1}{10} = (x + \ln(x))|_{1}^{10}\]